Ticket Number: HEA-NOC/20020606-01 Ticket Status: CLOSE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> Janet
Ticket Owner: [log in to unmask] Problem Fixer: NTL
Ticket Opened: Problem Start: 20020606 14:19 UTC+1
Ticket Closed: 20020617 12:33:57 UTC Problem End: 20020613 17:36:32 UTC+1
BER alarms on link between HEAnet and JANET.
all traffic with janet rerouted via geant links
20020606 1430 CW Janet also confirm link down, they have been doing no work on thier end
20020606 1432 CW Report fault to NTL, awaiting ticket ref #
20020606 1534 CW NTL ref # 140171
20020606 1625 CW NTL report fault is at the belfast end, refer them to janet operations to arrange site access
20020606 1644 CW NTL unable to go onsite until tomorrow, circuit has come up at 20020606 1635 and counters have been cleared
20020607 1030 cw NTL yet to go onsite, fault reoccoured at 1002
20020607 1552 cw NTL engineer Hugh in cardif reports that a faulty line card was swaped yesterday, resilience will be restored early next week. NTL will continue to investigate BER alarms.
20020611 11:39 AH Contact NTL for update. They report that the circuit is still running on the backup resilient link and that works will be scheduled this week for repair and return to the main path. The BER alarms on 7/6/02 have been reported as related to work being carried out by an engineer onsite in Belfast at the time.
20020611 15:30 NTL report that the primary path is testing clear and that they are preparing to restore resiliance.
20020613 12:10 cw NTL report that work to restore resiliance this evening after 17:30 hrs and that work should take not more than 30 minutes
20020614 NTL confirm that work is complete. We will continue to monitor.
20020617 No further errors. Ticket closed.
Time to Fix:
2 hours 30 mins